Greta has 45 minutes before her Job interview. She is 15 miles away from the interview. How fast must she drive to be on time?
thanx

Answer by stanbon(48516)

Greta has 45 minutes before her Job interview. She is 15 miles away from the interview. How fast must she drive to be on time?
You want an answer in mph or miles/hr.
So convert 45 min to 3/4 hr.
Then consider 15miles/(3/4)hr.= (60/3)mi/hr = 20mph
Cheers,
Stan H.